QUESTION 22

  1. A firm can produce 450 gallons of milk per day with 4 workers and 500 gallons per day with 5 workers. The marginal product of the fifth worker expressed in gallons per worker per day, is:

    35.

    50.

    70.

    350.

2.5 points   

QUESTION 23

  1. Exhibit 7-1 Production of pizza data
    Workers Pizzas
    0 0
    1 4
    2 10
    3 15
    4 18
    5 19


    Exhibit 7-1 shows the change in the production of pizzas as more workers are hired. The marginal product of the second employee equals:

    4.

    10.

    14.

    6.

    15.

2.5 points   

QUESTION 24

  1. Exhibit 7-1 Production of pizza data
    Workers Pizzas
    0 0
    1 4
    2 10
    3 15
    4 18
    5 19


    Exhibit 7-1 shows the change in the production of pizzas as more workers are hired. The marginal product of the fifth employee equals:

    4.

    18.

    19.

    3.

    1.

2.5 points   

QUESTION 25

  1. Exhibit 7-8 Costs schedules for producing pizza

    Pizzas
    Fixed
    Cost
    Variable
    Cost
    Total
    Cost
    Marginal
    Cost
    0 $      $    $      $   
    1     5
    2 13
    3 10
    4 100 140
    5 20
    6 85
    7 215


    By filling in the blanks in Exhibit 7-8, the marginal cost of the sixth pizza is shown to be equal to:

    $10.

    $15.

    $20.

    $25.

    $85.

2.5 points   

QUESTION 26

  1. Exhibit 7-12 Cost schedule for producing pizza

    Pizzas
    Fixed
    Cost
    Variable
    Cost
    Total
    Cost
    0 $    $    $   
    1 48
    2 17
    3 27
    4 78
    5 40
    6 64
    7 80


    By filling in the blanks in Exhibit 7-12, the AFC of 4 pizzas is shown to be equal to:

    $10.

    $9.50.

    $19.50.

    $40.

    $78.

2.5 points   

QUESTION 27

  1. Exhibit 7-12 Cost schedule for producing pizza

    Pizzas
    Fixed
    Cost
    Variable
    Cost
    Total
    Cost
    0 $    $    $   
    1 48
    2 17
    3 27
    4 78
    5 40
    6 64
    7 80


    By filling in the blanks in Exhibit 7-12, the AVC of 4 pizzas is shown to be equal to:

    $10.

    $9.50.

    $19.50.

    $40.

    $78.

2.5 points   

QUESTION 28

  1. Exhibit 7-12 Cost schedule for producing pizza

    Pizzas
    Fixed
    Cost
    Variable
    Cost
    Total
    Cost
    0 $    $    $   
    1 48
    2 17
    3 27
    4 78
    5 40
    6 64
    7 80


    By filling in the blanks in Exhibit 7-12, the ATC of 4 pizzas is shown to be equal to:

    $10.

    $9.50.

    $19.50.

    $40.

    $78.
Answer #1

Ans.22- 50

Marginal product of fifth worker= 500-450= 50

Ans.23- 6 (10-4 = 6)

Ans.24- 1 (19-18=1)

Ans.25- 25

Ans.26- 10

AFC = 40/4 = 10.